OpenStudy (anonymous):

HELP! What is the mole fraction of KCl in a 0.20 molal solution of KCl? (molar mass of KCl = 75 g; molar mass of water = 18 g)

OpenStudy (anonymous):

@thomaster

thomaster (thomaster):

0.20 molar, that means 0.2 mol per liter How many moles water are in 1 liter? (1 gram water = 1 ml)

OpenStudy (anonymous):

Ml stands for moles?

OpenStudy (anonymous):

H2O molar mass = 18.0152 g/mol 1000 g / 18.0152 g/mol = 55.51 mol????

thomaster (thomaster):

That's correct. Now we have 0.2 mol KCl in 55.51 mol H2O Do you know how to calculate the KCl mole fraction?

OpenStudy (anonymous):

no

thomaster (thomaster):

\(\sf\large mole\ fration = \dfrac{moles\ solute}{total\ moles\ in\ solution}\)

OpenStudy (anonymous):

ok 0.2/55.51?

thomaster (thomaster):

almost right, the total moles in solution is everything added up. so 55.51+0.2 That makes \(\sf\dfrac{0.2}{55.71}\)

OpenStudy (anonymous):

oh okay. that would be 0.00359002??

thomaster (thomaster):

Yes :)

OpenStudy (anonymous):

that's it??

OpenStudy (anonymous):

can you help me with some questions? please! I only have 9 left

thomaster (thomaster):

Sure, make sure to close this question and post a new one for every question.